OpenStudy (anonymous):

-y + 3y - 6y + 12y solve?

HanAkoSolo (jamierox4ev3r):

first add the like terms together @ashleyjade

HanAkoSolo (jamierox4ev3r):

think of it as all the terms that match go together. For example -y+3y=2y. In this problem, all of the terms are alike. If that is the case, then what do you get?

HanAkoSolo (jamierox4ev3r):

heeellooo? XD

OpenStudy (anonymous):

i.e. compute \[-1+3-6+12\] and stick a \(y\) next to it

HanAkoSolo (jamierox4ev3r):

right as all the terms have the variable y, they are all like terms, and if you add them all together, that would be your answer. Just don't forget to stick the y-variable to your number as satellite said

OpenStudy (anonymous):

at the end i just stick y next to the #?

OpenStudy (anonymous):

-1 +9?

HanAkoSolo (jamierox4ev3r):

what satelite said is that you add all of the numbers together, but we only took off the variable to show you that these are like terms

OpenStudy (anonymous):

yes, stick the \(y\) next to it

OpenStudy (anonymous):

answer would be 2y?

HanAkoSolo (jamierox4ev3r):

so: -y+3y-6y+12y really equals -1+3-6+12 because the terms are all like. THen, you put the variable next to the end of the above(^) because there is a y t9 begin with so it must stay

OpenStudy (anonymous):

\[-1\times \clubsuit+3\times \clubsuit-6\times \clubsuit+12\times \clubsuit\] \[=(-1+3-6+12)\times \clubsuit\] \[=8\times \clubsuit\]

HanAkoSolo (jamierox4ev3r):

Idk if this helps, but just think of this problem as a simplification problem ;)

HanAkoSolo (jamierox4ev3r):

lol @satellite73 are those clubs supposed to represent the y's? O_O Anyhow, lets see if this helps you @ashleyjade : -y+3y=2y -1+3=2 Using this little model, can you figure the rest out?